OpenStudy (anonymous):

solve for x 1/x + 1/x+16 = 1/15

OpenStudy (saifoo.khan):

\[\frac 1x + \frac{1}{x+16} = \frac{1}{15}\]that?

OpenStudy (anonymous):

answer is X= 480

OpenStudy (saifoo.khan):

that can be marked as wrong. as word-case matters in math. :O

OpenStudy (anonymous):

i think it is \[x=-10,x=24\]

OpenStudy (anonymous):

yes, the equation you wrote is the right one saifoo

OpenStudy (saifoo.khan):

Sat is right.

OpenStudy (anonymous):

yup, you're right saifoo, thank you!

OpenStudy (saifoo.khan):

No probs! (:

OpenStudy (anonymous):

2/x +1/16=1/15 2/x=1/15-1/16 2/x=1/240 1/x=1/ 480 , so x= 480

OpenStudy (saifoo.khan):

anetika, check your equation!! :o

OpenStudy (anonymous):

thanks satellite!

OpenStudy (anonymous):

then you get \[\frac{2x+16}{x(x+16)}=\frac{1}{15}\]

OpenStudy (anonymous):

yw

OpenStudy (saifoo.khan):

why tht shows me im typing when im not even? >.<

OpenStudy (anonymous):

ohh , i was wrong.. thank u saifoo

OpenStudy (saifoo.khan):

No probs.